# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ Acknowledgments
We would like to thank Dmitry Gayfulin, who contributed to the development of the regularization technique at an early stage of this study. The helpful comments of two anonymous reviewers are highly appreciated. The study has been conducted under Task 1.1.1 of the Scientific Research and Technology Plan of the Russian Federal Service for Hydrometeorology and Environmental Monitoring.
Appendix A Linear sequential filtering
Here we review the Kalman filter and sub-optimal non-ensemble and ensemble filters that are used in our numerical experiments. For more details and a broader perspective, see, e.g., Asch et al. (2016), Houtekamer and Zhang (2016).
@@ A.1 Linear filtering setup
The goal of sequential filtering is to estimate the current state of a system having (i) a forecast model capable of predicting the future state of the system given its current state, (ii) a set of observations on the system, and (iii) an observation model, which relates the observations to the system state (the truth).
In the time-discrete filtering, observations are available at discrete time instants. The estimates of the truth are sought at the same time instants. A sequential filter consists of alternating short-term forecast (time update) and analysis (observation update) steps.
In the time-discrete linear filtering, the evolution of the truth ${\bf x}_{t}$ (a vector of length $n_{\rm x}$) between consecutive analyses is assumed to be governed by the stochastic equation
${\bf x}_{t}={\bf F}_{t}{\bf x}_{t-1}+\boldsymbol{\varepsilon}_{t},$ (24)
where ${\bf F}_{t}$ is the known forecast-model operator (a matrix) and $\boldsymbol{\varepsilon}_{t}$ is the unknown model error assumed to be a random vector with mean zero and known covariance matrix ${\bf Q}_{t}$. The assimilated observations ${\bf x}^{\rm obs}_{t}$ (a vector of length $n_{t}^{\rm obs}$) are assumed to be linearly related to the truth:
${\bf x}^{\rm obs}_{t}={\bf H}_{t}{\bf x}_{t}+\boldsymbol{\eta}_{t},$ (25)
where ${\bf H}_{t}$ is the known observation-model operator and $\boldsymbol{\eta}$ is the unknown random observation error vector, which has zero mean and known covariance matrix ${\bf R}$.
The time-discrete linear filtering aims at finding mean-square optimal estimates and their uncertainties. A mean-square optimal estimate is a linear combination of all available data (observations) such that the estimation error is orthogonal (with respect to the covariance inner product) to the data.
@@ A.2 Forecast step
In the optimal (Kalman) filter, having the optimal analysis vector ${\bf x}^{\rm a}_{t-1}$ (the point estimate of ${\bf x}_{t-1}$ from observations ${\bf x}^{\rm obs}_{t-1},{\bf x}^{\rm obs}_{t-2},\dots$) implies that the optimal forecast vector (the point estimate of ${\bf x}_{t}$ from the same observations) is simply
@@
provided that $\boldsymbol{\varepsilon}_{t}$ is uncorrelated with the observation error $\boldsymbol{\eta}_{t^{\prime}}$ for any $t^{\prime}<t$. Its error covariance matrix is easily seen to be ${\bf B}_{t}={\bf F}_{t}{\bf A}_{t-1}{\bf F}_{t}^{\textsf{T}}+{\bf Q}_{t}$, where ${\bf A}$ stands for the analysis-error covariance matrix. So, the Kalman filter propagates in time the optimal point estimate of the truth and its error covariance matrix.
In a sub-optimal (ensemble or non-ensemble) filter, the forecast Eq. (26) is no longer optimal but it is unbiased if so is ${\bf x}^{\rm a}_{t-1}$ (subtract Eq. (24) from Eq. (26)). For this reason, we use Eq. (26) to compute the control forecast in the experiments presented in this paper (see also the remark in A.4).
Instead of propagating the error covariance matrix, the ensemble filter propagates an ensemble that represents the uncertainty in ${\bf x}^{\rm f}_{t}$:
${\bf x}^{{\rm fe}(\mu)}_{t}={\bf F}_{t}{\bf x}^{{\rm ae}(\mu)}_{t-1}+\boldsymbol{\varepsilon}_{t}^{(\mu)},$ (27)
where $\mu=1,2,\dots,M$ labels the ensemble member, $M$ is the ensemble size, the analysis-ensemble members ${\bf x}^{{\rm ae}(\mu)}_{t-1}$ are defined below in A.4, and the simulated model error $\boldsymbol{\varepsilon}_{t}^{(\mu)}$ is drawn from a probability distribution that, ideally (and in the experiments in this study), coincides with the distribution of the true model error $\boldsymbol{\varepsilon}_{t}$.
@@ A.3 Non-ensemble analysis
In a non-ensemble analysis scheme, we are given, first, the forecast (analysis background) ${\bf x}^{\rm f}$, see Eq. (26). (Till the end of this section, the time subscripts are omitted since all vectors correspond to the same time instant.) Second, we have current observations ${\bf x}^{\rm obs}$. Third, a static proxy ${\bf B}_{\rm static}$ to the true background-error covariance matrix ${\bf B}$ is available. The point estimate of the truth (often called the deterministic or control analysis) ${\bf x}^{\rm a}$ is sought as a linear combination of ${\bf x}^{\rm f}$ and ${\bf x}^{\rm obs}$.
If (i) the true background-error covariance matrix ${\bf B}$ is known, (ii) ${\bf x}^{\rm f}$ is an optimal estimate of the truth by itself (meaning that no affine transformation of the forecast is better in the mean-square sense), and (iii) $\boldsymbol{\eta}$ is probabilistically independent of the forecast error $\boldsymbol{\xi}={\bf x}^{\rm f}-{\bf x}$ (i.e., of all past observation and model errors), then the optimal (Kalman filter) analysis scheme can be employed:
@@
In a sub-optimal non-ensemble analysis scheme, only a proxy, ${\bf B}_{\rm static}$, to the true background-error covariance matrix is available. In this situation, the common practice is to plug in ${\bf B}_{\rm static}$ instead of ${\bf B}$ into Eq. (29) and use the resulting gain matrix to compute the analysis vector using Eq. (28). This approach is taken in so-called variational data assimilation, e.g., Asch et al. (2016).
@@ A.4 Ensemble analysis
An ensemble analysis scheme has on input the control forecast ${\bf x}^{\rm f}$ and the forecast ensemble $\{{\bf x}^{{\rm fe}(\mu)}\}_{\mu=1}^{M}$, see Eq. (27). The static covariance matrix ${\bf B}_{\rm static}$ can also be available.
An ensemble analysis scheme computes the control analysis vector ${\bf x}^{\rm a}$ and the analysis ensemble $\{{\bf x}^{{\rm ae}(\mu)}\}_{\mu=1}^{M}$ aimed to characterize the error in ${\bf x}^{\rm a}$.
@@
Remark. With nonlinear forecast models, the control forecast ${\bf x}^{\rm f}$ is normally replaced with the ensemble mean forecast $\overline{{\bf x}^{\rm fe}}$. However, with the linear forecast model, linear observations, and zero-mean forecast and observation errors, the control forecast is unbiased (see below in this section and A.2), and is, therefore, a better proxy to the truth than the ensemble mean. Indeed, $\overline{{\bf x}^{\rm fe}}-{\bf x}\equiv(\overline{{\bf x}^{\rm fe}}-{\bf x}^{\rm f})+({\bf x}^{\rm f}-{\bf x})=-\frac{1}{M}\sum\boldsymbol{\xi}^{(\mu)}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}^{(\mu)}={\bf x}^{\rm fe}-{\bf x}^{{\rm fe}(\mu)}$ are the pseudo-random ensemble perturbations and $\boldsymbol{\xi}$ is the error in the control forecast. If $\operatorname{\mathbb{E}}\boldsymbol{\xi}=0$ and $\boldsymbol{\xi}^{(\mu)}$ are independent of $\boldsymbol{\xi}$, then the mean square error of the ensemble mean is seen to be greater than that of the control forecast. For this reason, we use ${\bf x}^{\rm f}$ rather than $\overline{{\bf x}^{\rm fe}}$ to compute the deterministic analysis in our experiments.
@@
The ensemble analysis computes the gain matrix, $\widehat{\bf K}$, using Eq. (29) in which ${\bf B}$ is replaced with an estimate of $\widehat{\bf B}$ computed from the ensemble ${\bf x}^{{\rm fe}(\mu)}$ and, possibly, from static covariances ${\bf B}_{\rm static}$. Recent-past estimated covariances can also be utilized, see references in Introduction, item 4. The Ensemble Kalman Filter (EnKF) explicitly or implicitly relies on the sample covariance matrix ${\bf S}$ to specify the prior covariances. In its so-called stochastic variant (which is used in our experiments), $\widehat{\bf B}={\bf B}_{\rm EnKF}$ is specified as ${\bf S}$ element-wise multiplied by a so-called localization matrix. Hybrid ensemble-variational data assimilation schemes use a convex combination of ${\bf B}_{\rm static}$ and ${\bf B}_{\rm EnKF}$ to specify $\widehat{\bf B}$.
@@
The resulting gain matrix $\widehat{\bf K}$ is used to compute the deterministic analysis ${\bf x}^{\rm a}$ following Eq. (28). Since $\widehat{\bf K}$ is sub-optimal, ${\bf x}^{\rm a}$ is not optimal either, but it is unbiased if so is the control forecast. Indeed, writing $\widehat{\bf K}=\widehat{\bf K}(\boldsymbol{\xi}^{(1)},\boldsymbol{\xi}^{(2)},\dots,\boldsymbol{\xi}^{(M)})$ to indicate its dependence on the forecast ensemble perturbations $\boldsymbol{\xi}^{(\mu)}$, substituting $\widehat{\bf K}$ into Eq. (28), and using Eq. (25), we obtain the analysis error
${\bf x}^{\rm a}-{\bf x}=\boldsymbol{\xi}+\widehat{\bf K}(\boldsymbol{\xi}^{(1)},\boldsymbol{\xi}^{(2)},\dots,\boldsymbol{\xi}^{(M)})\cdot(\boldsymbol{\eta}-{\bf H}\boldsymbol{\xi}),$ (30)
@@
which has zero expectation because $\boldsymbol{\xi}$ and $\boldsymbol{\eta}$ are zero-mean random vectors independent of $\boldsymbol{\xi}^{(\mu)}$.
@@
${\bf x}^{{\rm ae}(\mu)}={\bf x}^{{\rm fe}(\mu)}+\widehat{\bf K}\cdot({\bf x}^{\rm obs}+\boldsymbol{\eta}^{(\mu)}-{\bf H}{\bf x}^{{\rm fe}(\mu)}),$ (31)
where $\boldsymbol{\eta}^{(\mu)}$are independent pseudo-random draws from the true observation-error distribution.
Appendix B Spherical harmonics, Fourier-Legendre transform, and white noise
The normalization of the spherical harmonics is such that they are orthonormal with respect to the inner product $(\psi_{1},\psi_{2})=\int_{{\mathbb{S}}^{2}}\psi_{1}(x)\psi_{2}^{*}(x)\,{\rm d}x$, where $\psi_{1}$ and $\psi_{2}$ are square integrable functions on the sphere and the asterisk denotes complex conjugation
The normalization of the Legendre polynomials is such that $P_{\ell}(1)=1$.
The addition theorem of spherical harmonics reads
where $x$ and $y$ are two points on the unit sphere and $\rho(x,y)$ is the great-circle distance between them.
The forward Fourier-Legendre transform of the function $F(\rho)$ (where $\rho\in[0,\pi]$) is the set of real numbers $\widetilde{F}_{\ell}$ (where $\ell=0,1,\dots,\ell_{\rm max}$) defined as
$\widetilde{F}_{\ell}=2\pi\int_{0}^{\pi}F(\rho)P_{\ell}(\cos\rho)\sin\rho\,{\rm d}\rho.$ (33)
$\widetilde{F}_{\ell}$ are seen to be the coefficients of the expansion of $F(\rho)$ in Legendre polynomials:
$F(\rho)=\sum^{\ell_{\rm max}}_{\ell=0}\frac{2\ell+1}{4\pi}\,\widetilde{F}_{\ell}\,P_{\ell}(\cos\rho).$ (34)
We call Eq. (34) the backward Fourier-Legendre transform.
By standard Gaussian white noise, we mean a generalized random process, $\alpha(x)$, that is, a random continuous linear functional that acts on a non-random test function $\phi(x)$ returning a zero-mean Gaussian random variable denoted as $\int\phi(x)\,\alpha(x)\,{\rm d}x$ whose variance is $\int|\phi(x)|^{2}\,{\rm d}x$. In the spectral expansion of the white noise,
$\alpha(y)=\sum^{\infty}_{\ell=0}\sum^{\ell}_{m=-\ell}\widetilde{\alpha}_{\ell m}\,Y_{\ell m}(y),$ (35)
$\widetilde{\alpha}_{\ell m}$ are mutually uncorrelated complex-valued random Fourier coefficients with $\operatorname{\mathbb{E}}\widetilde{\alpha}_{\ell m}=0$ and ${\mathsf{Var}}\,\widetilde{\alpha}_{\ell m}=1$. More specifically, $\widetilde{\alpha}_{l0}$ are real-valued and all the other $\widetilde{\alpha}_{\ell m}$ are complex circularly symmetric random variables (e.g., Gallager, 2013, section 3.7) such that $\widetilde{\alpha}_{l,-m}=\widetilde{\alpha}_{\ell m}^{*}$. The band-limited white noise is defined by Eq. (35), in which summation over $\ell$ is truncated at $\ell_{\rm max}<\infty$.
Appendix C Non-negative definiteness of the non-stationary model
@@
Appendix D The spatial model on the circle
@@
where the real-valued convolution kernel $u(x,z)$ is an even function of the second argument. Let us employ the truncated Fourier series expansion of the kernel in the form
$u(x,z)=\frac{1}{\sqrt{2\pi}}\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\sigma_{\ell}(x)\,\mathrm{e}^{\mathsf{i}\ell z}.$ (40)
Here $\sigma_{\ell}(x)$ are real-valued functions such that $\sigma_{-\ell}(x)=\sigma_{\ell}(x)$. With the positive-definite kernel assumption (section 2.4), $\sigma_{\ell}(x)>0$.
In the spectral expansion of the band-limited Gaussian white noise,
@@
$B(x,x^{\prime})=\operatorname{\mathbb{E}}\xi(x)\,\xi(x^{\prime})=\sum\,\sigma_{\ell}(x)\,\sigma_{\ell}(x^{\prime})\,\mathrm{e}^{\mathsf{i}\ell(x^{\prime}-x)}.$ (43)
Appendix E The most localized kernel in the stationary case
Here we consider a stationary process convolution model. We show that its convolution kernel is not unique given the model’s covariance function (or, equivalently, the spectrum). Uniqueness can be achieved by requiring that the kernel be most spatially localized.

### v5

@@ Acknowledgments
We would like to thank Dmitry Gayfulin, who contributed to the development of the regularization technique at an early stage of this study. Constructive criticism of two anonymous reviewers of the first (rejected) version of the manuscript is highly appreciated. Valuable comments of two anonymous reviewers and the editor, Alfred Stein, of the second version of the manuscript allowed us to substantially improve and clarify the manuscript. The study has been conducted under Task 1.1.1 of the Scientific Research and Technology Plan of the Russian Federal Service for Hydrometeorology and Environmental Monitoring.
Appendix A Approximate sequential filtering
We review the Kalman filter and sub-optimal non-ensemble and ensemble filters that are used in our numerical experiments. The sub-optimal (approximate) filters are reviewed in the linear setting sufficient for our purposes in this paper. For more details and a broader perspective, see, e.g., Asch et al. (2016), Houtekamer and Zhang (2016).
@@ A.1 Linear filtering setup
Here, we add a few notational and technical details to what is presented in section 2 above.
By $n_{\rm x}$ we denote the dimensionality of the system, i.e., the length of its state vector ${\bf x}_{t}$. By $n_{t}^{\rm obs}$ we denote the number of observations available at the time instant $t$. The model-error covariance matrix is $\operatorname{\mathbb{E}}\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t}^{\textsf{T}}={\bf Q}_{t}$. The observation-error covariance matrix is $\operatorname{\mathbb{E}}\boldsymbol{\varepsilon}_{t}\boldsymbol{\varepsilon}_{t}^{\textsf{T}}={\bf R}_{t}$.
The time-discrete linear filtering aims at finding mean-square optimal estimates and their uncertainties. A mean-square optimal estimate is a linear combination of all available data (observations) such that the estimation error is orthogonal (with respect to the covariance inner product) to the available data.
@@ A.2 Forecast step
In the optimal (Kalman) filter, assume we have the optimal analysis vector ${\bf x}^{\rm a}_{t-1}$, that is, the point estimate of ${\bf x}_{t-1}$ whose error is orthogonal to the observations ${\bf x}^{\rm obs}_{t-1},{\bf x}^{\rm obs}_{t-2},\dots$, along with the analysis error covariance metrix ${\bf A}_{t-1}$. Then, the optimal forecast vector (the point estimate of ${\bf x}_{t}$ from the same observations) is simply the deterministic (control) forecast
@@
provided that the model error $\boldsymbol{\eta}_{t}$ is uncorrelated with the observation error $\boldsymbol{\varepsilon}_{t^{\prime}}$ for any $t^{\prime}<t$. It is straightforward to write down the error covariance matrix of the deterministic forecast: ${\bf B}_{t}={\bf F}_{t}{\bf A}_{t-1}{\bf F}_{t}^{\textsf{T}}+{\bf Q}_{t}$. So, the Kalman filter propagates in time the optimal point estimate of the truth and its error covariance matrix.
In a sub-optimal (ensemble or non-ensemble) filter, the forecast Eq. (38) is no longer optimal but it is unbiased if so is ${\bf x}^{\rm a}_{t-1}$ (subtract Eq. (1) from Eq. (38)). For this reason, we use Eq. (38) to compute the control forecast in the experiments presented in this paper (see also Remark in A.4).
Instead of propagating the error covariance matrix, an ensemble filter propagates an ensemble that represents the uncertainty in ${\bf x}^{\rm f}_{t}$:
${\bf x}^{{\rm fe}(k)}_{t}={\bf F}_{t}{\bf x}^{{\rm ae}(k)}_{t-1}+\boldsymbol{\eta}_{t}^{(k)},$ (39)
where $k=1,2,\dots,K$ labels the ensemble member, $K$ is the ensemble size, the analysis-ensemble members ${\bf x}^{{\rm ae}(k)}_{t-1}$ are defined below in A.4, and the simulated model error $\boldsymbol{\eta}_{t}^{(k)}$ is drawn from a probability distribution that, ideally (and in the experiments in this study), coincides with the distribution of the true model error $\boldsymbol{\eta}_{t}$.
@@ A.3 Non-ensemble analysis
In a non-ensemble analysis scheme, we are given, first, the forecast (analysis background) ${\bf x}^{\rm f}$, see Eq. (38). (Till the end of this subsection, the time subscripts are omitted since all vectors correspond to the same time instant.) Second, we have current observations ${\bf x}^{\rm obs}$. Third, a static proxy ${\bf B}_{\rm static}$ to the true background-error covariance matrix ${\bf B}$ is available. The point estimate of the truth (often called the deterministic or control analysis) ${\bf x}^{\rm a}$ is sought as a linear combination of ${\bf x}^{\rm f}$ and ${\bf x}^{\rm obs}$.
If (i) the true background-error covariance matrix ${\bf B}$ is known, (ii) ${\bf x}^{\rm f}$ is an optimal estimate of the truth by itself (meaning that no affine transformation of the forecast is better in the mean-square sense), and (iii) $\boldsymbol{\varepsilon}$ is probabilistically independent of the forecast error $\boldsymbol{\xi}={\bf x}^{\rm f}-{\bf x}$ (this is the case if $\boldsymbol{\varepsilon}$ is independent of all past observation and model errors), then the optimal (Kalman filter) analysis scheme can be employed:
@@
In a sub-optimal non-ensemble analysis scheme, only a proxy ${\bf B}_{\rm static}$ to the true background-error covariance matrix is available. In this situation, the common practice is to plug in ${\bf B}_{\rm static}$ instead of ${\bf B}$ into Eq. (41) and use the resulting gain matrix to compute the analysis vector using Eq. (40). This approach is taken in so-called variational data assimilation (e.g. Asch et al., 2016).
@@ A.4 Ensemble analysis
The ensemble analysis scheme has on input the control forecast ${\bf x}^{\rm f}$ and the forecast ensemble $\{{\bf x}^{{\rm fe}(k)}\}_{k=1}^{K}$, see Eq. (39). The static covariance matrix ${\bf B}_{\rm static}$ can also be available. The goal is to compute the deterministic analysis vector ${\bf x}^{\rm a}$ and the analysis ensemble $\{{\bf x}^{{\rm ae}(k)}\}_{k=1}^{K}$ aimed to characterize the error in ${\bf x}^{\rm a}$.
@@
Remark. With nonlinear forecast models, the deterministic forecast ${\bf x}^{\rm f}$ is normally replaced with the ensemble mean forecast $\overline{{\bf x}^{\rm fe}}$. However, with the linear forecast model, linear observations, and zero-mean forecast and observation errors, the deterministic forecast is unbiased (see below in this subsection and A.2). Therefore it is a better proxy to the truth than the ensemble mean. Indeed, $\overline{{\bf x}^{\rm fe}}-{\bf x}\equiv(\overline{{\bf x}^{\rm fe}}-{\bf x}^{\rm f})+({\bf x}^{\rm f}-{\bf x})=-\frac{1}{K}\sum\boldsymbol{\xi}^{(k)}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}^{(k)}={\bf x}^{\rm fe}-{\bf x}^{{\rm fe}(k)}$ are the pseudo-random ensemble perturbations and $\boldsymbol{\xi}$ is the error in the control forecast. If $\operatorname{\mathbb{E}}\boldsymbol{\xi}=0$ and $\boldsymbol{\xi}^{(k)}$ are independent of $\boldsymbol{\xi}$, then the mean square error of the ensemble mean is seen to be greater than that of the control forecast. For this reason, we used ${\bf x}^{\rm f}$ rather than $\overline{{\bf x}^{\rm fe}}$ to compute the deterministic analysis in our experiments.
@@
The ensemble analysis computes the gain matrix $\widehat{\bf K}$ using Eq. (41) in which ${\bf B}$ is replaced with its estimate, $\widehat{\bf B}$, computed from the ensemble ${\bf x}^{{\rm fe}(k)}$ and, possibly, from the static covariances ${\bf B}_{\rm static}$. Recent-past estimated covariances can also be utilized, see references in section 1.1, item 4. The Ensemble Kalman Filter (EnKF) explicitly or implicitly relies on the sample covariance matrix ${\bf S}$ to specify the prior covariances. In the so-called stochastic variant of EnKF (which is used in our experiments), $\widehat{\bf B}$ equals ${\bf B}_{\rm EnKF}$, which is ${\bf S}$ element-wise multiplied by a so-called localization matrix. Hybrid ensemble-variational data assimilation schemes use a convex combination of ${\bf B}_{\rm static}$ and ${\bf B}_{\rm EnKF}$ to specify $\widehat{\bf B}$.
@@
The resulting gain matrix $\widehat{\bf K}$ is used to compute the deterministic analysis ${\bf x}^{\rm a}$ following Eq. (40). Since $\widehat{\bf K}$ is sub-optimal, ${\bf x}^{\rm a}$ is not optimal either, but it is unbiased if so is the control forecast. Indeed, writing $\widehat{\bf K}=\widehat{\bf K}(\boldsymbol{\xi}^{(1)},\boldsymbol{\xi}^{(2)},\dots,\boldsymbol{\xi}^{(K)})$ to indicate its dependence on the forecast ensemble perturbations $\boldsymbol{\xi}^{(k)}$, substituting $\widehat{\bf K}$ into Eq. (40), and using Eq. (2), we obtain the analysis error
${\bf x}^{\rm a}-{\bf x}=\boldsymbol{\xi}+\widehat{\bf K}(\boldsymbol{\xi}^{(1)},\boldsymbol{\xi}^{(2)},\dots,\boldsymbol{\xi}^{(K)})\cdot(\boldsymbol{\varepsilon}-{\bf H}\boldsymbol{\xi}),$ (42)
@@
which has zero expectation because $\boldsymbol{\xi}$ and $\boldsymbol{\varepsilon}$ are zero-mean random vectors independent of $\boldsymbol{\xi}^{(k)}$.
@@
${\bf x}^{{\rm ae}(k)}={\bf x}^{{\rm fe}(k)}+\widehat{\bf K}\cdot({\bf x}^{\rm obs}+\boldsymbol{\varepsilon}^{(k)}-{\bf H}{\bf x}^{{\rm fe}(k)}),$ (43)
where $\boldsymbol{\varepsilon}^{(k)}$ are independent pseudo-random draws from the true observation-error distribution.
Appendix B Non-negative definiteness of the non-stationary model
@@
Appendix C The spatial model on the circle
@@
where the real-valued convolution kernel $u(x,\rho)$ is an even function of the second argument. Let us employ the truncated Fourier series expansion of the kernel in the form
$u(x,\rho)=\frac{1}{\sqrt{2\pi}}\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\sigma_{\ell}(x)\,\mathrm{e}^{\mathsf{i}\ell\rho}.$ (48)
Here $\sigma_{\ell}(x)$ are real-valued functions such that $\sigma_{-\ell}(x)=\sigma_{\ell}(x)$. With the positive-definite convolution kernel (section 3.4), $\sigma_{\ell}(x)>0$. The band-limited Gaussian white noise $\alpha$ in Eq. (47) has the spectral expansion
@@
$B(x,x^{\prime})=\operatorname{\mathbb{E}}\xi(x)\,\xi(x^{\prime})=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\,\sigma_{\ell}(x)\,\sigma_{\ell}(x^{\prime})\,\mathrm{e}^{\mathsf{i}\ell(x^{\prime}-x)}.$ (51)
Appendix D The most localized kernel in the stationary case
Here we consider the stationary process convolution model. We show that its convolution kernel is not unique given the model’s covariance function (or, equivalently, the spectrum). Uniqueness can be achieved by requiring that the kernel be most spatially localized.